Please help me solve this. It is a Mathematical Expression. Liam has z pennies. How many dollars does he have? Write an expression that models this situation.
Answer by jim_thompson5910(35256)

100 pennies = 1 dollar

1 penny = 1/100 dollar

z pennies = z/100 dollars


So if he has z pennies, then he has z/100 dollars.


Ex: he has z = 300 pennies, so he has z/100 = 300/100 = 3 dollars.